Determine the magnitude and direction of the following forces in a 150N box that lies on a 30 degree inclined plane by whose coeffient of kinetic friction is 0.32


1
Expert's answer
2021-05-30T13:27:52-0400

"N=F\cos\alpha=150\cdot 0.87=130~N,"

"F_{fr}=\mu N=0.32\cdot 130=42~N."